Ingraham Inc. currently has $800,000 in accounts receivable, and its days sales outstanding (DSO) is 65 days. It wants to reduce its DSO to 30 days by pressuring more of its customers to pay their bills on time. If this policy is adopted, the company's average sales will fall by 25%. What will be the level of accounts receivable following the change? Assume a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Credit sales = Accounts receivables *365/Days sales outstanding

= 800,000*365/65

=$4,492,307.69

Revised sales = 4,492,307.69*75% =$3,369,230.77

Accounts receivables after change = credit sales*desired DSO/365

=$276,923.08